# ESG-Net: Event-Aware Semantic Guided Network for Dense Audio-Visual Event Localization

### v1

@@ IV-C Ablation Study
To analyze the impact of different components on our model in detail, we conduct ablation experiments based on features extracted by ONE-PEACE [13]. Note that we keep two decimal places in this section for thorough comparison.
@@
Key Components in ESI. As shown in Table III, we present the impact of each component on the performance within the our ESI module. Firstly, as shown in the first three rows of Table III, we remove $\mathcal{L}_{AV1}$, $\mathcal{L}_{AV2}$, and $\mathcal{L}_{AV3}$, respectively, to analyze the role of different component of multi-stage semantic guidance. It can be seen that $\mathcal{L}_{AV3}$ has the most significant impact on performance, while $\mathcal{L}_{AV1}$ and $\mathcal{L}_{AV2}$ have some impact on performance. This indicates that by employing multi-stage semantic guidance for cross-modal semantic bridging at multiple intermediate layers, the capability of our model to localize audio-visual events can be enhanced.
Secondly, the table shows the results of sequentially removing $\mathcal{L}_{AV3}$, $\mathcal{L}_{AV2}$, and $\mathcal{L}_{AV1}$ in the third to fifth rows. As the number of stages with semantic guidance decreases, the performance drops accordingly. Note that there is a special case here, where the average mAP when $\mathcal{L}_{mcls}=\alpha_{1}\mathcal{L}_{AV1}$ (52.49%) is higher than when $\mathcal{L}_{mcls}=\alpha_{1}\mathcal{L}_{AV1}+\alpha_{2}\mathcal{L}_{AV2}$ (51.82%). We guess this might be due to the early multi-modal representations $\hat{F_{g}}$ in Audio(Visual)-Driven Mixture can introduce much noise without sufficient cross-modal consistency constraint. Our multi-stage semantic guidance can not only achieve semantic bridging between the two branches but also indirectly optimize the early multi-modal representation $\hat{F_{g}}$ by reducing the differences in audio and video features.
@@
Finally, the last four rows of the table demonstrate the contribution of $\mathcal{L}_{mcls}$ and the early fusion branch to our model: (1) When only the multi-modal early fusion is used, there is almost no improvement (only 0.01%) on the performance compared to the original backbone [1]; (2) When only the multi-stage semantic guidance ($\mathcal{L}_{mcls}$) is used, the model’s performance shows a significant improvement; (3) When both the multi-modal early fusion and $\mathcal{L}_{mcls}$ are used, the model achieves the highest average mAP, and the results at the tIoU thresholds of [0.5:0.1:0.9] are also the best or the second-best. These phenomena indicate that the performance improvement brought by the multi-modal early fusion is not due to an increase in parameters. In fact, it is the collaboration of the multi-modal early fusion with the multi-stage semantic guidance that enables better event-related semantic bridging in intermediate layers.
(a) DAVE in a quiet environment.
(b) DAVE in a noisy environment.
@@
Figure 3: Qualitative results on the testing set. "GT" means ground truth, "Base" denotes the baseline model as described in Section IV-A. We present the audio-visual event detection results and the predictions of the temporal boundaries for each event. We add extension lines to all boundaries of our model for clear comparison.
Impact of the Weights in $\mathcal{L}_{mcls}$. As in Eq. (8), $\mathcal{L}_{mcls}$ is the weighted sum of $\mathcal{L}_{AV1}$, $\mathcal{L}_{AV2}$, and $\mathcal{L}_{AV3}$, the results in Table IV show how the weights of our multi-stage semantic guidance loss functions affects the final performance. In this part, we present all the results at the tIoU thresholds [0.1:0.1:0.9] for clear comparison. By comparing the first two rows of Table IV, even the semantic guidance between modalities is stronger when $\alpha_{i}=0.5$ $(i=1,2,3)$, the average mAP is approximately equal when $\alpha_{1}=0.3$, $\alpha_{2}=0.4$, $\alpha_{3}=0.5$. This indicates that gradually enhancing semantic guidance in three stages can effectively achieve cross-modal semantic bridging, thereby improving the detection capability of the model. Next, by gradually increasing the weight difference $(\alpha_{i}-\alpha_{i-1})$ between adjacent stages, we can observe that larger difference can better improve the effectiveness of semantic guidance. It can be seen that when $\alpha_{1}=0.3$, $\alpha_{2}=0.6$, $\alpha_{3}=0.9$, the mAP is either the best or the second-best at low tIoU thresholds (from 0.1 to 0.6). A similar phenomenon can also be observed by comparing the second and third rows of Table IV. This indicates that as the weight difference increases, the explicit semantic guidance can enhance the detection capability of our model, but may obscure the its understanding of the temporal boundaries of audio-visual events. We guess this may be because when the classification loss of the later layer $\mathcal{L}_{AV3}$ is increased, the model pays more attention on event classification, thereby relatively weakening the final generalized IoU loss $\mathcal{L}_{reg}$ used for regressing the temporal boundaries of events.
@@
Impact of the Number of Experts in Each MoDE Layer. We study the impact of the number of experts in each layer of MoDE on the performance, and the results are shown in Table V. It can be seen that the mAP reaches the highest value of 54.74% when the number of experts is set to 2. And the detection accuracy of the model is either the best or the second-best at the tIoU thresholds of [0.5:0.1:0.9]. However, as the number of experts increases in each layer, the performance of our model gradually declines. Because only one expert can be activated in each layer during one training iteration, adding too many experts in each layer may lead to insufficient training for each expert.
@@
Impact of the MoE Layers in MoDE. Table VI shows how the number of MoE layers in MoDE affects the performance of the model. When the number of MoE layers is 16, mAP reaches the best at 54.87%, and when the number of MoE layers is 4, mAP achieves the second-best at 54.74%. Different from the conclusion in Table V, although increasing the number of MoE layers also increases the number of experts, the proportion of experts activated during one training iteration remains constant. Increasing the number of MoE layers can achieve more complex combinations of experts as described in Section III-D (each MoE layer has $n$ experts, $m$ MoE layers would have $n^{m}$ possible combinations), which may improve the final performance. However, the training and inference time increases linearly with the number of MoE layers. 16-layer MoE improves mAP by 0.13% over 4-layer MoE, but the inference time increases by 20.04%. Therefore, after comprehensive consideration of the two metrics, we ultimately adopted the 4-layer MoE which is more efficient.
@@
TABLE VII: Comparison of parameters and computational load.
Method Parameters(M) FLOPs(G) Avg.
UnAV [1] 140.79 60.28 51.0
UniAV(STF) [5] 186.00 32.83 51.7
LoCo [6] 103.73 31.45 53.4
Ours 86.13 20.36 54.7
IV-D Parameter and Computational Load.
@@
To demonstrate the effectiveness and efficiency of our method, we compare the parameter and FLOPs of different methods as shown in Table VII. Note that we only compare state-of-the-art DAVE methods that are either open-source or well-documented in papers. It can be seen that our method not only achieves the highest mAP of 54.7% but also has the lowest consumption in terms of memory usage and computational load. In fact, the main computational consumption of our method is concentrated on ESI, as this module is mainly used for complex temporal modeling and fusion of single-modal and multi-modal features. While MoDE has only about 2M parameters, but it can efficiently model dependencies among multiple events. In summary, our ESG-Net is both effective and efficient, which enables it to better focus on event-related content with lower computational costs.
IV-E Visualization and Discussion
@@
Qualitative Comparison. We present the qualitative detection results of our method in Figure 3 and compare with the baseline method [1] ("Base"). Due to the multiple detection results for the each audio-visual event, we only select the result with the maximum tIoU with "GT" while ensuring that every event in "GT" has at least one corresponding detection result. To better show the performance of ESG-Net and "Base", we add extension lines for each detected event. As we can see, in quiet environments like Figure 3 (a), for the events "baby babbling", "baby laughter" and "people shouting", both our method and "Base" are capable of accurately detecting the categories of these events. However, the temporal boundaries regressed by "Base" are often either too early or too late. As for the events "kid speaking", although our model can not accurately output the three sets of temporal boundaries, it still achieved a higher tIoU, especially in the noisy background between 22 seconds and 35 seconds. In Figure 3 (b), for the events "ambulance siren" and "police car siren", both our method and "Base" show significant deviations of temporal boundaries, which are due to the strong interference in noisy environments. But for the events "vehicle honking", "Base" has multiple missed detections, while our method can still accurately detect these events, which demonstrates the robustness of ESG-Net. Overall, these results show the superior performance of detection and boundaries regression of our method, even in noisy environments.
@@
Impact of Multi-Stage Semantic Guidance. We show the role of our multi-stage semantic guidance by the attention weight matrixs across three stages (Single-Modal Attention, Audio(Visual)-Driven Mixture, Cross-Modal Pyramid). There are four examples in Figure 4 to show the effectiveness of the multi-stage semantic guidance. The first and second rows of each example are the results of the visual branch and the audio branch, respectively. Each red bounding box denotes a audio-visual event. As we can see in example (a)(b)(c), in stage 1, the model does not have a clear focus area, showing lower concern for audio-video events. Additionally, the attention weight matrixs have significant differences between audio and visual branch. In stage 2, our model begins to pay attention to the event-related areas, but their attention to local details remained insufficient. While in stage 3, both the audio branch and the visual branch can basically pay attention to audio-visual events, and the attention weight matrices of the two branches become similar. There is a special case in example (d), where the video contains only one event. From stage 1 to stage 3, the visual branch can well locate the event. However, the audio branch in stage 1 occurs false detections from 0 to 25 snippets and from 170 to 210 snippets. But that noise is gradually filtered out through cross-modal semantic guidance in subsequent stages. To sum up, the above results demonstrate that our multi-stage semantic guidance can progressively guide different modalities to focus on event-related information while gradually filtering out irrelevant noise, thereby improving the final detection performance.
@@ V Conclusion
In this paper, we propose an event-aware semantic guided network that includes an early semantics interaction module and a mixture of dependency experts module for dense audio-visual event localization. ESI introduces an early fusion branch and implements multi-modal consistency constraints through multi-stage semantic guidance, which achieves cross-modal bridging in intermediate layers and allows model to focus on event-related content. MoDE reveals the multi-event dependencies by adaptively combining suitable experts across different MoE layers. In summary, ESG-Net performs explicit semantic alignment in intermediate layers and multi-event relationships extraction, achieving more accurate detection and localization for audio-visual events. Through detailed ablation experiments, we demonstrate the significant role of each component in ESG-Net. Our method not only markedly outperforms state-of-the-art methods in terms of accuracy, but also shows efficient performance in parameters and computational load. In the future, we plan to incorporate advanced pre-trained knowledge to further enhance the capability of our model to process and understand audio-visual information.

### v2

@@ IV-D Ablation Study
To analyze the impact of different components on our model in detail, we conduct ablation experiments based on features extracted by ONE-PEACE [17].
@@
Key Components in ESI. As shown in Table IV, we present the impact of each component on the performance within our ESI module. Firstly, as shown in the first three rows of Table IV, we remove $\mathcal{L}_{AV1}$, $\mathcal{L}_{AV2}$, and $\mathcal{L}_{AV3}$, respectively, to analyze the role of different component of multi-stage semantic guidance. It can be seen that $\mathcal{L}_{AV3}$ has the most significant impact on performance, while $\mathcal{L}_{AV1}$ and $\mathcal{L}_{AV2}$ have some impact on performance. The results from the fourth to sixth rows in this table exhibit the same pattern. These results indicate that by employing multi-stage semantic guidance for cross-modal semantic bridging, the capability of our model to localize audio-visual events can be enhanced.
@@
Secondly, the results from the fourth to sixth rows in Table IV individually demonstrate the contribution of $\mathcal{L}_{AVi}$ in ESI, while the last four rows demonstrate the contribution of $\mathcal{L}_{mcls}$ and the early fusion branch to our model: (1)Without early multi-modal representations, $\mathcal{L}_{AVi}$ can bring some improvement for event detection, but the performance remains limited by the feature diffierence between two single-modal branches; (2) When only the early multi-modal representations are used, there is almost no improvement on the performance compared to the original backbone [1]; (3) When only the multi-stage semantic guidance ($\mathcal{L}_{mcls}$) is used, the model’s performance shows a significant improvement; (4) When both the early multi-modal representations and $\mathcal{L}_{mcls}$ are used, the model achieves the highest average mAP, and the results at the tIoU thresholds of [0.5:0.1:0.9] are also the best or the second-best. These phenomena indicate that the performance improvement brought by the early multi-modal representations is not due to an increase in parameters. In fact, it is the collaboration of the multi-modal early fusion with the multi-stage semantic guidance that enables better event-related semantic bridging in intermediate layers.
Finally, we replace the ESI module with alternative architectures, the results are presented in Table V. Here, "Cat" denotes using simple concatenation fusion at the early fusion stage; "CMPTE" refers to replacing ESI with the Cross-Modal Pyramid Transformer Encoder from UnAV [1]; and "CMCC" indicates substituting ESI with the Cross-Modal Consistency Collaboration module from CCNet [38]. It can be seen that our method achieves significantly superior performance compared to alternative module configurations, with an average improvement of 3.87% (±0.4%) while maintaining comparable parameters with these methods.
@@
Figure 3: Qualitative results on the testing set. "GT" means ground truth, "Base" denotes the baseline model.
Impact of different structures in MoDE. We replace the MoDE module with alternative structures, the results are presented in Table VI. "MBM" denotes using C*C-matrix-based modeling for event dependencies; "GCN" refers to replacing MoDE with multiple cascaded graph convolutional networks; "Attention" refers to modeling event dependencies by self-attention; "MoS2 Block" refers to adding transformer block after each layer of MoDE and applying expert routing like [74] (we fail to add the shift modules due to dimension mismatch); "TDM" indicates replacing MoDE with the Temporal Dependency Modeling from UnAV [1]. Our method achieves the best or the second-best at the tIoU thresholds of [0.5:0.1:0.9] compared with the above structures. Although the average mAP of "TDM" is close to our method, its parameter count exceeds that of MoDE by an order of magnitude. The "MoS2 Block" also adopts a MoE architecture for video processing, but the lack of pre-trained weight results in interference between transformer layers and MoE layers. Finally, compared to static architectures like ’MBM’ and ’GCN’, MoDE can adaptively uncover relationships among co-occurrring events through dynamic expert selection, demonstrating superior performance.
@@
Impact of the Number of Experts in Each MoDE Layer. We study the impact of the number of experts in each layer of MoDE on the performance, and the results are shown in Table VII. It can be seen that the mAP reaches the highest value of 55.1% when the number of experts is set to 2. And the detection accuracy of the model is always the best at the tIoU thresholds of [0.5:0.1:0.9]. However, as the number of experts increases in each layer, the performance of our model gradually declines. Because only one expert can be activated in each layer during one training iteration, adding too many experts in each layer may lead to insufficient training for each expert.
@@
Impact of the MoE Layers in MoDE. Table VIII shows how the number of MoE layers in MoDE affects the performance. When the number of MoE layers is 16, mAP reaches the second-best at 54.9%, and when the number of MoE layers is 4, mAP achieves the best at 55.1%. It can be seen that although increasing the number of MoE layers also increases the total number of experts, the proportion of experts activated during one training iteration remains constant. Increasing the number of MoE layers can achieve more complex combinations of experts as described in Section III-D (each MoE layer has $n$ experts, $m$ MoE layers would have $n^{m}$ possible combinations), which may improve the final performance. However, the training and inference time grows linearly as the number of MoE layers increases. The inference time of 16-layer MoE increases by 20.04% over 4-layer MoE. After comprehensive consideration of the two metrics, we ultimately adopted the 4-layer MoE which is both effective and efficient.
@@
Figure 5: Qualitative results of average performance on co-occurring events with different overlap rates on the videos from the test set.
TABLE IX: Comparison of parameters and computational load.
Method Params.(M) FLOPs(G) Mem.(M) Tr. Time(h) Avg.
UnAV[1] 140.79 60.28 3628 16.03 51.0
UniAV[5] 186.00 32.83 4304 9.04 51.7
LoCo[6] 103.73 31.45 - - 53.4
CCNet[38] 238.78 72.14 5790 20.31 54.1
Ours 86.13 20.36 2617 22.50 55.1
IV-E Parameter and Computational Load.
@@
To demonstrate the effectiveness and efficiency of our method, we compare the parameter, FLOPs, training time and memory usage of different methods as shown in Table IX. Note that we only compare state-of-the-art DAVE methods that are either open-source or well-documented in their papers. It can be seen that our method not only achieves the highest mAP of 55.1% but also has the lowest consumption in terms of parameter, FLOPs and memory usage. In fact, the main computational consumption of our method is concentrated on ESI, as this module is mainly used for complex temporal modeling and fusion of single-modal and multi-modal features. While MoDE has only about 2M parameters, but it can efficiently model dependencies among multiple events. Secondly, our method exhibits the longest training duration among all compared approaches, which is mainly attributed to the sequential structure within the MoDE module that introduces extended forward propagation pathways. However, the long training duration of CCNet [38] is caused by the scale of its model and high computational demands. In summary, our ESG-Net achieves the best comprehensive performance, significantly reducing computational load while maintaining superior focus on event-related content.
IV-F Visualization and Discussion
@@
Qualitative Comparison. We present the qualitative detection results of our method in Figure 3 and compare with the baseline method [1] ("Base"). To better show the performance of ESG-Net and "Base", we add extension lines for each detected event. As we can see, for the events "baby babbling", "baby laughter" and "people shouting", both our method and "Base" are capable of accurately detecting the categories of these events. However, the temporal boundaries regressed by "Base" are often either too early or too late. As for the events "kid speaking", although our model can not accurately output the three sets of temporal boundaries, it still achieved a higher tIoU, especially in the noisy background between 22 seconds and 35 seconds. Overall, these results show the superior performance of detection and boundaries regression of our method.
@@
Impact of Multi-Stage Semantic Guidance. We show the role of our multi-stage semantic guidance by the attention weight matrixs across three stages (Single-Modal Attention, Audio(Visual)-Driven Mixture, Cross-Modal Pyramid) in Figure 4. The first and second rows of each example are the results of the visual branch and the audio branch, respectively. Each red bounding box denotes a audio-visual event. As we can see, in stage 1, the model does not have a clear focus area, showing lower concern for audio-video events. Additionally, the attention weight matrixs have significant differences between audio and visual branch. In stage 2, our model begins to pay attention to the event-related areas, but their attention to local details remained insufficient. While in stage 3, both the audio branch and the visual branch can basically pay attention to audio-visual events, and the attention weight matrices of the two branches become similar. To sum up, the above results demonstrate that our multi-stage semantic guidance can progressively guide different modalities to focus on event-related information while gradually filtering out irrelevant noise.
Detection capability for co-occurring events. To evaluate the effectiveness of our MoDE in detecting co-occurring events, we test the mAP on videos that contained co-occurring events with different overlap rates. "UnAV_woTDM" denotes the UnAV [1] without Temporal Dependency Modeling, and "Ours_woMoDE" represents ESG-Net without MoDE. We conduct experiments using features extracted by I3D+VGGish [15, 16], with the results shown in Figure 5. We observed that compared to models that do not consider inter-event modeling ("ActionFormer" [32] and "UnAV_woTDM"), both UnAV [1] and our method demonstrate more stable performance when handling the videos with any overlap rates. Compared to UnAV [1], our method achieves better performance on the videos with overlap rates of (0, 0.5] and (0.8, 1.0], demonstrating the effectiveness of MoDE in capturing event dependencies. However, a slight performance degradation of our method is observed on videos with overlap rates in (0.5, 0.8]. This can be attributed to the limited number of test samples within this specific interval, which may lead to bias in the evaluation metrics.
@@ V Conclusion
In this paper, we propose an event-aware semantic guided network that includes an early semantic interaction module and a mixture of dependency experts module for dense audio-visual event localization. ESI introduces the multi-stage feature fusion and implements multi-modal consistency constraints through multi-stage semantic guidance, which achieves cross-modal bridging in intermediate layers and allows model to focus on event-related content. MoDE reveals the multi-event dependencies by adaptively combining suitable experts across different MoE layers. In summary, ESG-Net performs explicit semantic alignment in intermediate layers and multi-event relationships extraction, achieving more accurate detection and localization for audio-visual events. Through detailed ablation experiments, we demonstrate the individual contribution of each component in ESG-Net. Our method not only markedly outperforms state-of-the-art methods in terms of accuracy, but also shows efficient performance in parameters and computational load.
We end with some limitations and outlining directions for future research. First, we plan to incorporate our method with adaptive temporal attention mechanism and uncertainty-aware modeling to address noisy constraints in single branches. Secondly, current DAVE methods are limited to fixed event categories and lacks detailed textual descriptions of video events. We propose to leverage the pre-trained knowledge in multi-modal large language models by fine-tuning with mixed question-answering data, to develop a unified framework capable of handling temporal-aware audio-visual downstream tasks (e.g., dense audio-visual event localization, video captioning and audio-visual video parsing) in open scenarios.